OpenStudy (anonymous):

Please help! I'm stuck on this question: Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (−6, 3).

OpenStudy (anonymous):

easy man,, see 1st we know that its gradient is 3/2 as its perpendicular to the line y = 3/2x + 4 now we have to find y-intercept 3=(3/2)*-6 + y-intercept 3+9=y-intercept y-intercept= 12 There equation of line will be y=3/2x + 12

OpenStudy (anonymous):